Find the HCF of 75 and 243 and express it in the form of 75x + 243y . 

Grade:10

1 Answers

Debarati
2 years ago
75= 31*52= 3*5*5
243= 35= 3*3*3*3*3
HCF= 31= 3
LCM= 35*52= 3*3*3*3*3*5*5= 6075
\thereforeThe HCF of 75 and 243 is 3.

I hope you understood.